7 A DOUBLE CHANNEL 2 -CRXP E ) Ad A TENNIO MEMBER HOW MUCH PARA LAS CAN BE APPLIED I ts 3k Fy-50 TYPE (1972) F4336 Hoooo C

A Double Channel 2-C8x18.75 is used as a tension member. How much factored load can be applied?

Both Part A and B need to be answered for the question. Type A, (A572) Channel = 50ksi, (A36) Gusset PL = 36ksi

Type B (A36) Channel & Gusset PL = 36ksi

Thank you!
Answer #1

Cijven chamnd member: - 2 (8 x 18.75 - Oo91 Grade :- A36 member is connected with the web a poop; Given channel section to 24. Fracture in the Net section in An= Agon(dent = 1102- 14) 11 (0.487) 9.072 in 2 Connection hength in Loading direction LaguGross section of Tension Agt = ax l length (6-2)]xt 214)(0.487) Net = 3.896 in? section of tension Ant = 2x [length (b-c)-In-d Design strength of bolt in shear, Ørns (0.175):10.5) (Fub) mA = 10:75).10.5) (120) 14) 10.6013) - 54.12 kips & Design StreI Gross section .. of Tension Agt = [hength (b-1)]xt = (4)(0.375) = 1.5 in z Ant = [hength (6-1)-(n-1) doh] H) --[4-(1911]0-31 fiz 58 Isi : Type B stul groole. A 36. . For Az6 steel, - ty = 36 ksi, tu= 58 ksi, bott dia do= 2 elia of bott hole dpAny = 4 [hength la-b) – (m-1)+0:5 )ek} lt) = 4111- 13.5) 10.87597 *0.487 - 14.488 in2 Similarly, Agr - 214010487) - 3.896 ini Design shear strength / bolt = 10.75) (015) (150) (4) (0.4498) = 49-7 kips | Design bearing strength=${24 fude] =1075) (2.4=195-9 kips : Design strength of Connection Min{®, ® ©} =195-9 kips ... Maximum factored hood Pu=195.9kAll 3 checks are performed.